OpenStudy (adrianna.gongora):

Will Medal and Fan ;-; Just Help me Regan is trying to find the equation of a quadratic that has a focus of (–2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (adrianna.gongora):

@naate

OpenStudy (anonymous):

this is not nearly has hard as it looks what is halfway between the horizontal line \(y=13\) and the point \((-2,5)\)?

OpenStudy (anonymous):

which is the same question as "what is halfway between 5 and 13?" or even "what is the average of 5 and 13?"

OpenStudy (adrianna.gongora):

9

OpenStudy (anonymous):

i agree, so the vertex is this sucker is \((-2,9)\)

OpenStudy (anonymous):

do you know whether this faces up or down?

OpenStudy (adrianna.gongora):

um down?

OpenStudy (anonymous):

ummm yes the directrix is above the focus, so it opens down

OpenStudy (anonymous):

equation will be \[(y-k)^2=4p(x-h)\] you know \(x=-2,k=5\) all that is left is \(p\)

OpenStudy (adrianna.gongora):

so how do I get p?

OpenStudy (anonymous):

to find it, \(p\) is the distance between the focus and the vertex, i.e. between \((-2,5)\) and \((-2,9)\) or equivalently the distance between the vertex \((-2,9)\) and the ditrectrix \(y=13\)

OpenStudy (anonymous):

how far from 5 to 9?

OpenStudy (adrianna.gongora):

4

OpenStudy (adrianna.gongora):

so it's (y−k)2=4p(x−h) x=−2,k=5 p=4

OpenStudy (anonymous):

yes, but make sure to put a minus sign in front of the \(4p\)since it opens down let me know what you get

OpenStudy (adrianna.gongora):

so it's (y−k)2=4p(x−h) x=−2,k=5 -4p

OpenStudy (anonymous):

silly rabbit, replace the variables by the numbers !

OpenStudy (adrianna.gongora):

;-; ohhhh so it's so it's (y−5)2=4(4)(-2−h)

OpenStudy (anonymous):

\[(y−k)^2=4p(x−h)\] replace \(h\) by \(-2\) replace \(k\) by \(9\) and replace \(p\) by \(-4\)

OpenStudy (anonymous):

\(k=9\) not \(5\)

OpenStudy (anonymous):

and you need to replace \(h\) by \(-2\), not \(x\)

OpenStudy (adrianna.gongora):

\[(y−9)=4(4)(-2−h)\]

OpenStudy (anonymous):

closer, not yet

OpenStudy (adrianna.gongora):

\[(y−5)2=4(4)(x−-2)\]

OpenStudy (anonymous):

not \(-2-h\) put \(h=-2\) not \(x\) that is your variable

OpenStudy (anonymous):

closer still, but \(p=-4\) because it opens down also if you write \(--2\) your teacher will think you are daft just make it \(+2\)

OpenStudy (adrianna.gongora):

\[(y−5)2=4(-4)(x+2)\] did I do it right this time?

OpenStudy (anonymous):

what would be even better is \[(y-9)^2=-16(x+2)\]

OpenStudy (anonymous):

don't forget vertex was \((-2,9)\)

OpenStudy (anonymous):

want to check it?

OpenStudy (anonymous):

oh damn i made a huge mistake, sorry

OpenStudy (adrianna.gongora):

djcnpasijdb ijdcijkdc j ;-; that is my mind right now I dislike math right now ;-; and yes it would be better I forgot all about the vertex ;-;

OpenStudy (anonymous):

good thing i checked the answer

OpenStudy (anonymous):

since it opens down the x term is squared, not the y term try \[-16(y-9)=(x+2)^2\]

OpenStudy (anonymous):

my mistake, sorry

OpenStudy (adrianna.gongora):

is that the answer?

OpenStudy (adrianna.gongora):

@satellite73

OpenStudy (adrianna.gongora):

@clayton4christ ?

OpenStudy (anonymous):

Sorry. I am not at all good at math

OpenStudy (adrianna.gongora):

;-; its okay